# Optimized Memory Tagging on AmpereOne® Processors

### v2

@@ VIII Discussion and Future Work
Co-location of tags with data by using a small number of ECC bits for tag storage has significant advantages for a datacenter SoC - no memory capacity impact and low performance overhead for a broad range of datacenter class workloads. However, as Sections IV and VII point out, there are impacts to enabling MTE in the platform hardware, even if tag checking is not being used by software. For the AmpereOne® SoC’s ECC tag implementation, the impact is a small (-1% to -3%) performance delta due to the need to perform Read-Modify-Write operations when preserving tags for full-cacheline stores that would otherwise be Write-Only. Memory controller designs can consider optimizing this flow further to reduce total overhead. Also, as stated in Section VI, the memory controller implementation in the AmpereOne® SoC for option D provides 100% of the fault detection, 100% correction of bounded faults, and 99.98% correction of unbounded faults. As outlined in [ecc2022] the "level of 9’s" correction of unbounded faults can be increased with additional support in the memory controller implementation. Such improvements can be achieved by carrying error information forward from one codeword to one or more subsequent codewords in the same cacheline. By using this technique, which depends on the extreme unlikelihood of errors occurring in different DRAM devices in the same cacheline, the level of 9’s can be improved significantly, to as high as 99.9999%.
@@
The AmpereOne® processor uses bits reserved in the mesh interconnect for metadata to transport tag values alongside the data. The ARM CHI-E extension added explicit support for tag management and transport in the mesh protocol [AMBA_CHI_Spec]. Future Ampere devices that support the CHI-E mesh protocol will leverage the tag management support for transporting tag values with data using these protocol extensions.
Industry standards such as the Compute Express Link (CXL-3.1) provide profiles for CXL devices with memory, referred to as CXL.mem or CXL Type 3 configurations [CXLSpec]. Enhancements to the specification for carrying metadata have been defined in the CXL 3.2 specification [CXLSpec3_2]. Future MTE implementations leverage these extensions to support memory tagging for the CXL.Mem/Type 3 devices. The support requires commercial availability of memory technology devices that support the CXL 3.2 metadata extension along with sufficient storage for tags in the metadata bits.
@@ IX Conclusion
X Acknowledgements
This work would not have been possible without the contributions of many of our colleagues, including the designers and verification engineers that implemented and validated MTE. We would like to specifically thank our performance analysis team: Olivier Singla, Jiangning Liu, Feng Xue, James Burke, Sean Daniels, and Steve Winiecki for conducting extensive performance measurements, and Syed Fakhri for his support of the performance monitoring infrastructure. We are also grateful to Richard Shannon, Mark Charney, Shahid Khan, and Shravan Narayan for their valuable guidance and review of this paper.

### v3

@@ VIII Discussion and Future Work
Co-location of tags with data by using a small number of ECC bits for tag storage has significant advantages for a datacenter SoC - no memory capacity impact and low performance overhead for a broad range of datacenter class workloads. However, as Sections IV and VII point out, there are impacts to enabling MTE in the platform hardware, even if tag checking is not being used by software. For the AmpereOne® SoC’s ECC tag implementation, the small performance delta is due to the need to perform Read-Modify-Write operations when preserving tags for full-cacheline stores that would otherwise be Write-Only. Memory controller designs can consider optimizing this flow further to reduce total overhead. Also, as stated in Section VI, the memory controller implementation in the AmpereOne® SoC for option D provides 100% of the fault detection, 100% correction of bounded faults, and 99.98% correction of unbounded faults. As outlined in [69] the "level of 9’s" correction of unbounded faults can be increased with additional support in the memory controller implementation. Such improvements can be achieved by carrying error information forward from one codeword to one or more subsequent codewords in the same cacheline. By using this technique, which depends on the extreme unlikelihood of errors occurring in different DRAM devices in the same cacheline, the level of 9’s can be improved significantly, to as high as 99.9999%.
@@
The AmpereOne® processor uses bits reserved in the mesh interconnect for metadata to transport tag values alongside the data. The ARM CHI-E extension added explicit support for tag management and transport in the mesh protocol [8]. Future Ampere devices that support the CHI-E mesh protocol will leverage the tag management support for transporting tag values with data using these protocol extensions.
Industry standards such as the Compute Express Link (CXL-3.1) provide profiles for CXL devices with memory, referred to as CXL.mem or CXL Type 3 configurations [20]. Enhancements to the specification for carrying metadata have been defined in the CXL 3.2 specification [23]. Future MTE implementations leverage these extensions to support memory tagging for the CXL.Mem/Type 3 devices. The support requires commercial availability of memory technology devices that support the CXL 3.2 metadata extension along with sufficient storage for tags in the metadata bits.
@@ IX Conclusion
Acknowledgement
This work would not have been possible without the contributions of many of our colleagues, including the designers and verification engineers that implemented and validated MTE. We would like to specifically thank our performance analysis team: Olivier Singla, Jiangning Liu, Feng Xue, James Burke, Sean Daniels, and Steve Winiecki for conducting extensive performance measurements, and Syed Fakhri for his support of the performance monitoring infrastructure. We are also grateful to Richard Shannon, Mark Charney, Shahid Khan, and Shravan Narayan for their valuable guidance and reviews.